OpenStudy (ms_sp):

1. Kenny and Kyle are huge Colorado Avalanche fans. Kenny buys a T-shirt and two souvenir hockey pucks for $36, and Kyle spends $64 on two T-shirts and three pucks. Find the price of a T-shirt and the price of a puck. 2. Carol orders five White Castle hamburgers and small order of French for $4.44 and Omar orders four hamburgers and two small fries for $5.22. Find the cost of a hamburger and the cost of a small order of French fries at White Castle. 3. Through January 2014, Beyonce has been nominated for ten more bet awards than Jay-Z. Determine how many nominations, total of 50. will medal

OpenStudy (aveline):

Kenny buys one shirt and two hockey pucks. It totals to be $36. Let's make 's' represent shirts and 'p' represent pucks. s + 2p = $36 Kyle bought two shirts and three pucks and paid $64 2s + 3p = $64 Let's rewrite the first equation to isolate the variable 's': s + 2p - 2p = $36 - 2p s = $36 - 2p Now we know what 's' is equal to, so we can plug it into the second equation.

OpenStudy (ms_sp):

@Aveline what about others?

OpenStudy (aveline):

You use the exact same process for the other equations. For example, on number two, the two equations would be: 5h + f = $4.44 (where h=hamburgers and f=fries) 4h +2f = $5.22

OpenStudy (ms_sp):

I'm still confused! @Aveline